OpenStudy (anonymous):

Mr. Jones is asked to write two types of radical equations for the SAT Prep Course. The first solution to the radical equation must be extraneous. The second solution to the radical equation must be non-extraneous. Write one equation where the solution is extraneous. Then write a second equation where the solution is non-extraneous. Using complete sentences, explain each step when solving to justify your examples. Can you please help me out with this question?@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you familiar with the term "extraneous" ?

OpenStudy (anonymous):

isn't is when something is excluded from the original equation? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

in a way, yes

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

there's more to it though

jimthompson5910 (jim_thompson5910):

let's say we had this equation \[\Large \sqrt{x} = -1\] and we state that x is a real number

jimthompson5910 (jim_thompson5910):

what is the solution for that equation?

OpenStudy (anonymous):

I don't understand extraneous and non extraneous solutions very well, that is the main reason why i need help

jimthompson5910 (jim_thompson5910):

what operation undoes square roots?

OpenStudy (anonymous):

if you square something it undoes the square root.

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

square both sides of the equation I wrote. What do you get?

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

-1 times -1 is not -1

jimthompson5910 (jim_thompson5910):

(-1)^2 = (-1)*(-1) = ?

OpenStudy (anonymous):

sorry 1

jimthompson5910 (jim_thompson5910):

so x = 1 is a possible solution

jimthompson5910 (jim_thompson5910):

we need to check it

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{x} = -1\] \[\Large \sqrt{1} = -1\] \[\Large 1 = -1\] which is false

jimthompson5910 (jim_thompson5910):

so x = 1 is considered to be an "extraneous solution"

jimthompson5910 (jim_thompson5910):

it pops up in the process of solving, but it does NOT satisfy the original equation

jimthompson5910 (jim_thompson5910):

here is a page that has another example http://www.mathwords.com/e/extraneous_solution.htm

OpenStudy (anonymous):

but in the question it says that i also need a non-extraneous solution. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

did you check out the link I posted?

jimthompson5910 (jim_thompson5910):

it provides an example where 2 solutions pop up...but 1 of those solutions is extraneous

OpenStudy (anonymous):

which one is the non-extraneous?

jimthompson5910 (jim_thompson5910):

the one that makes the original equation true

jimthompson5910 (jim_thompson5910):

check both x = 9 and x = 4 with the original equation

OpenStudy (anonymous):

oh okay thanks!! is it okay if we go on twiddla later to check my work? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure we can do that

OpenStudy (anonymous):

thank you!! :)

jimthompson5910 (jim_thompson5910):

np

jimthompson5910 (jim_thompson5910):

let's find out

jimthompson5910 (jim_thompson5910):

so x is what value?

jimthompson5910 (jim_thompson5910):

that is the only possible solution for x

jimthompson5910 (jim_thompson5910):

you need to check it now

jimthompson5910 (jim_thompson5910):

so you are correct in stating that x = 81 is non-extraneous

jimthompson5910 (jim_thompson5910):

ok sounds good

OpenStudy (anonymous):

@jim_thompson5910 can u plz get on twiddla now?

OpenStudy (anonymous):

Lincoln and Gemma are looking at the equation the square root of the quantity of 2 times x minus 3 equals square root of x . Lincoln says that the solution is extraneous. Gemma says the solution is non-extraneous. Is Lincoln correct? Is Gemma correct? Are they both correct? Justify your response by solving this equation, explaining each step with complete sentences @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you have so far?

OpenStudy (anonymous):

nothing i need help with it

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

The equation is \[\Large \sqrt{2x}-3 = \sqrt{x}\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what happens when you square both sides

OpenStudy (anonymous):

idk i need help with it to understand

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{2x}-3 = \sqrt{x}\] \[\Large (\sqrt{2x}-3)^2 = (\sqrt{x})^2\] \[\Large (\sqrt{2x})^2-2*\sqrt{2x}*3+(3)^2 = (\sqrt{x})^2\] \[\Large 2x-6*\sqrt{2x}+9 = x\] \[\Large -6*\sqrt{2x}+9 = x-2x\] \[\Large -6*\sqrt{2x}+9 = -x\] \[\Large -6*\sqrt{2x} = -x-9\] \[\Large \sqrt{2x} = \frac{-x-9}{-6}\] \[\Large \sqrt{2x} = \frac{x+9}{6}\] I'll let you finish up

jimthompson5910 (jim_thompson5910):

sorry I made a typo, but I fixed it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm really confused :/ @jim_thompson5910 i'm so sorry

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (anonymous):

everything..

jimthompson5910 (jim_thompson5910):

you understand how I squared both sides right? I did so to undo the square root

jimthompson5910 (jim_thompson5910):

and I also used the rule (a-b)^2 = a^2 - 2ab + b^2 to expand out the left side

OpenStudy (anonymous):

okay then thats probably where i got stuck.. cuz idk that formula @jim_thompson5910

jimthompson5910 (jim_thompson5910):

that is the perfect square formula

jimthompson5910 (jim_thompson5910):

you would foil out (a-b)(a-b) to get a^2 - 2ab + b^2

jimthompson5910 (jim_thompson5910):

similar to (a+b)(a+b) = a^2 + 2ab + b^2

OpenStudy (anonymous):

okay so how would i answer the question?

jimthompson5910 (jim_thompson5910):

you would keep going to solve for x

OpenStudy (anonymous):

can u go on twiddla?

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{2x} = \frac{x+9}{6}\] \[\Large (\sqrt{2x})^2 = (\frac{x+9}{6})^2\] \[\Large 2x = \frac{(x+9)^2}{36}\]

jimthompson5910 (jim_thompson5910):

do you see how to solve that for x?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

(x+9)^2 turns into what expression

OpenStudy (anonymous):

(x+9)(x+9)?

jimthompson5910 (jim_thompson5910):

foil that out

OpenStudy (anonymous):

x^2+9x+9x+81?

jimthompson5910 (jim_thompson5910):

that turns into x^2 + 18x + 81

jimthompson5910 (jim_thompson5910):

\[\Large 2x = \frac{(x+9)^2}{36}\] \[\Large 2x = \frac{x^2+18x+81}{36}\] \[\Large 2x*36 = x^2+18x+81\] \[\Large 72x = x^2+18x+81\]

jimthompson5910 (jim_thompson5910):

the ultimate goal is to isolate x

OpenStudy (anonymous):

u keep adding stuff and that confuses me where did u get 36 from?

jimthompson5910 (jim_thompson5910):

6 square is 36, sry

OpenStudy (anonymous):

it's okay i'm just a bit slow math isn't my strongest subject

jimthompson5910 (jim_thompson5910):

\[\Large (\frac{x+9}{6})^2 = \frac{(x+9)^2}{6^2} \] \[\Large (\frac{x+9}{6})^2 = \frac{x^2+18x+81}{36} \]

jimthompson5910 (jim_thompson5910):

You're doing quite well actually

OpenStudy (anonymous):

not really.. hah i have no clue what's going on atm.

jimthompson5910 (jim_thompson5910):

well you do understand that squaring undoes square roots so that is usually the first thing to remember when it comes to problems like this

jimthompson5910 (jim_thompson5910):

the other stuff is just moving terms around, simplifying and such

jimthompson5910 (jim_thompson5910):

at first, all of this is very confusing, but with a lot of practice it should come easier

OpenStudy (anonymous):

:) thanks for the encouragement! really needed that :)

jimthompson5910 (jim_thompson5910):

You're welcome. You might find this hard to believe, but I also struggled with math too. I just kept at it and eventually things clicked. So don't give up.

OpenStudy (anonymous):

:) thank you so much!!

jimthompson5910 (jim_thompson5910):

see attached

jimthompson5910 (jim_thompson5910):

try this